5.One 55 g serving of a particular cereal supplies 270 mg of sodium, 11% of the recommended daily allowance. How many moles and atoms of sodium are in the recommended daily allowance?


1
Expert's answer
2022-06-08T02:05:04-0400

Solution:

270 mg of sodium – 11% of the recommended daily allowance

X mg of sodium – 100% of the recommended daily allowance

Using proportions to find out how many mg of sodium are in the recommended daily allowance:

X mg of sodium = (270 mg × 100%) / (11%) = 2454.545 mg

Therefore, the recommended daily allowance of sodium is 2454.545 mg


Convert mg to g:

(2454.545 mg) × (1 g / 1000 mg) = 2.454545 g = 2.455 g


The molar mass of sodium (Na) is 22.99 g/mol

Therefore,

Moles of Na = (2.455 g Na) × (1 mol Na / 22.99 g Na) = 0.10678 mol Na = 0.107 mol Na

The recommended daily allowance of sodium is 0.107 moles of Na


One mole of any substance contains 6.02×1023 atoms/molecules

Therefore,

Number of Na atoms = (0.107 mol) × (6.02×1023 atoms / 1 mol) = 6.44×1022 atoms

The recommended daily allowance of sodium is 6.44×1022 Na atoms


Answer: 0.107 moles and 6.44×1022 atoms of sodium are in the recommended daily allowance